|7-10|-|-5| =-2
(I don't understand and I can not work out how this equation works at all.

Answer by jim_thompson5910(35256)

First, 7-10 = -3, but remember that the absolute value of any number is positive. So |7 - 10| = |-3| = 3

This also applies to |-5| which means that |-5|=5

Put this all together to get:

|7 - 10| - | -5 | = | -3 | - | -5 | = 3 - 5 = -2


In the end, |7 - 10| - | -5 | = -2